# The deep multi-FBSDE method: a robust deep learning method for coupled FBSDEs

### v1

@@ 5 Conclusions and potential future research directions
In this paper, we introduced a two-phase deep-learning-based method, the deep multi-FBSDE method, for approximating coupled forward–backward stochastic differential equations (FBSDEs). Our approach targets scenarios where the standard deep BSDE method of Han, Jentzen, and E (2018) fails to converge. The main characteristic of our method is that its loss function is composed of the sum of multiple deep BSDE loss functions, each originating to a FBSDE satisfying the same partial differential equation (PDE) as the original FBSDE, but with transformed drift and driver. The method was demonstrated on two stochastic control problems and two FBSDEs related to advection diffusion reaction PDEs, unrelated to stochastic control. For neither of the problems, the deep BSDE method worked, but the deep multi-FBSDE method did. Overall, our results suggest that focusing on a family of equivalent FBSDEs offers a promising path to overcome convergence challenges in deep BSDE methods, thus broadening the scope of problems in finance, control theory, and related areas where these methods can be applied effectively.
Although the method seems to work very well on challenging FBSDE, we still do not understand the reason for this and why the deep BSDE method fails. There is thus a need to both understand theoretically for what problems the deep BSDE does not work and why, and why the deep multi-FBSDE method does work for these problems. From a practical perspective, the choice of $K$ and $\psi_{1},\dots,\psi_{K}$ in our algorithm is currently somewhat ad hoc. While this works for moderately complex problems, a more systematic approach to selecting $K$ and $\psi_{1},\dots,\psi_{K}$ would be especially beneficial when dealing with high-dimensional systems or models where it is difficult to anticipate the impact of different choices. Developing such a framework constitutes a natural next step toward enhancing both the reliability and the applicability of the proposed method. Furthermore, extending the methodology to broader classes of PDE-driven problems and exploring theoretical guarantees for large-scale or high-dimensional settings would deepen our understanding of its robustness and versatility.
@@ Acknowledgments
Part of this work was carried out during a two-month postdoctoral visit of Kristoffer Andersson (KA) at the University of Utrecht, whose hospitality and support are gratefully acknowledged. KA also acknowledges financial support from RiBa 2022 Research Fund provided by the University of Verona.

### v2

@@ 5 Conclusions and potential future research directions
In this paper, we introduced a two-phase deep-learning-based method, the deep multi-FBSDE method, for approximating coupled forward–backward stochastic differential equations (FBSDEs). Our approach targets scenarios where the standard deep BSDE method [17] fails to converge. The main characteristic of our method is that its loss function is composed of the sum of multiple deep BSDE loss functions, each originating to a FBSDE satisfying the same partial differential equation (PDE) as the original FBSDE, but with transformed drift and driver. The method was demonstrated on two stochastic control problems and two FBSDEs related to advection diffusion reaction PDEs, unrelated to stochastic control. For neither of the problems, the deep BSDE method worked, but the deep multi-FBSDE method did. Overall, our results suggest that focusing on a family of equivalent FBSDEs offers a promising path to overcome convergence challenges in deep BSDE methods, thus broadening the scope of problems in finance, control theory, and related areas where these methods can be applied effectively.
Although the method seems to work very well on challenging FBSDEs, we still do not understand the reason for this and why the deep BSDE method fails. There is thus a need to both understand theoretically for what problems the deep BSDE does not work and why, and why the deep multi-FBSDE method does work for these problems. From a practical perspective, the choice of $K$ and $\psi_{1},\dots,\psi_{K}$ in our algorithm is currently somewhat ad hoc. While this works for moderately complex problems, a more systematic approach to selecting $K$ and $\psi_{1},\dots,\psi_{K}$ would be especially beneficial when dealing with high-dimensional systems or models where it is difficult to anticipate the impact of different choices. Developing such a framework constitutes a natural next step toward enhancing both the reliability and the applicability of the proposed method. Furthermore, extending the methodology to broader classes of PDE-driven problems and exploring theoretical guarantees for large-scale or high-dimensional settings would deepen our understanding of its robustness and versatility.
@@ Acknowledgments
Part of this work was carried out during a two-month postdoctoral visit of Kristoffer Andersson (KA) at the University of Utrecht, whose hospitality and support are gratefully acknowledged. KA also acknowledges financial support from RiBa 2022 Research Fund provided by the University of Verona. We very much thank Per Ljung att Saab for his careful reading and many comments that helped us improve the manuscript, to Karl Hammar at Saab and Chalmers for pointing out an error, and to Benjamin Svedung Wettervik at Saab for some interesting reflections.